OpenStudy (anonymous):

A theater has 35 rows of chairs. The first row has 20 chairs, the second has 22, the third has 24, and so on. In total, the theater seats _____ people.

OpenStudy (saifoo.khan):

Arithmetic sequence?

OpenStudy (anonymous):

\[\sum_{i=0 \to 34}20+2{i}\]

OpenStudy (anonymous):

The sum of the first n terms of an arithmetic sequence is n/2[a1+an], where an=a1+(n-1)d, where d is the common difference, which is 2 in our case. Also n=35 and a35=20+34(2)=88, so S_35= 35/2[20+88]=1890